OpenStudy (anonymous):

Simplify 0.5(√K/√L) / 0.5(√L/√K).

OpenStudy (agreene):

\[\frac{0.5(\frac{\sqrt K}{\sqrt L})}{0.5 (\frac{\sqrt L}{\sqrt K})}\] Is that the question?

OpenStudy (anonymous):

Yep!

OpenStudy (agreene):

\[\frac{0.5 \sqrt {\frac{K}{L}}}{0.5 \sqrt{ \frac{L}{K}}}\] \[\sqrt{\frac{K^2}{L^2}}\] \[\frac{K}{L}\]